OpenStudy (anonymous):

I got the answer for x^2/4 + 4=0 but the answer was 4i, 4i what does that mean?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Imaginary_number

OpenStudy (zehanz):

i is a number with property i²=-1, or \(i=\sqrt{-1}\) If you solve \(\dfrac{x^2}{4}+4=0\), you get \(\dfrac{x^2}{4}=-4\), so \(x^2=-16\). Then \(x=\pm \sqrt{-16}=\pm \sqrt{16}\cdot\sqrt{-1}=\pm 4i\).

OpenStudy (zehanz):

If you are supposed to give answers like this (instead of: the given eqn. has no (real) solutions), I assume you have learned about these "imaginary numbers"...